OpenStudy (anonymous):

Solve Logarithm Equation . (a) log(x-2)-logx=-1

OpenStudy (anonymous):

I would first write the left as one log: log[(x-2)/x] = -1 (using the properties of logs) Then write it as an exponent: 10^-1 = (x-2)/x

OpenStudy (anonymous):

\[\log(x-2)/x=-1 \[(x-2)/x=e^-1\] \[(x-2)*e=x\] \[e*x-e*2=x\] e*x-x-x*2=0

OpenStudy (anonymous):

.1 = (x-2)/x .1x = x-2 -.9x=-2 x = 20/9

OpenStudy (anonymous):

its -1 t bates not +1 so how u write this 1 = (x-2)/x................ ???

OpenStudy (anonymous):

and\[\log _{e ^{(x-1)/x}} its the from of log

OpenStudy (anonymous):

THE ANSWER IS x=20/9

OpenStudy (anonymous):

logx/log2-logx=-1 logx-log2logx/log2=-1 logx-log2logx=-log2 logx(1-log2)=-log2 logx=-log2/(1-log2) x=2.22 =20/9

OpenStudy (anonymous):

10^(-1) = 0.1 which I used in my solution

OpenStudy (anonymous):

how is it ???its can't be cause \[\log _{e} \] or\[\log _{10}\] log base here u use 10 so u can write this problem as log_e and prosed

OpenStudy (anonymous):

anytime you see logx it's always assumed to be base 10. You could change the base to e but that would add extra steps

OpenStudy (anonymous):

but how can i learn that where i will use lag_e or log_10 sorry for my mistake sir T Bates i'm really sorry

OpenStudy (anonymous):

That's alright. Anytime you see log(x) it's base 10, ln(x) is base e.

OpenStudy (anonymous):

ok but when we use 2.303*ln with that time we shall write log why can u explain me sir

OpenStudy (anonymous):

I'm not sure what the question is... 2.303*ln (???) anytime you have a logarithm you need something inside the logarithm. But in this case this would be base e as ln is always base e